# Experiencer, Helper, or Observer: Online Fraud Intervention for Older Adults Through Role-based Simulation

### v1

@@ Abstract.
older adults, online fraud intervention, anti-fraud education, role-playing, simulation
copyright: acmcopyrightjournalyear: 2021doi: $X$.XXXX/XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 1journalnumber: 1article: 1publicationmonth: 1ccs: Security and privacy Usability in security and privacyccs: Human-centered computing Empirical studies in HCI
@@ 1. Introduction
Online fraud has emerged as one of the most prevalent and costly digital threats worldwide (Button et al., 2014; Cross, 2022). Older adults are disproportionately targeted due to multiple age-related risk factors (Judges et al., 2017; Alves and Wilson, 2008; James et al., 2014). Their vulnerability is further exacerbated by the continuous evolution and increasing sophistication of scam tactics (Deng et al., 2025; Zhai et al., 2025), as well as the use of AI to enhance the effectiveness, speed, and credibility of fraudulent schemes (Herrera et al., 2024). Consequently, older adults often express growing apprehension about the complexity of such threats and the difficulty of recognizing fraud in practice (Sheil et al., 2024). Fraud targeting older adults has increasingly become a global threat. In China, survey data from the Chinese Academy of Social Sciences revealed that 17.25% of older users had experienced online fraud in 2022 (Neican, 2022). According to the U.S. Federal Trade Commission, American victims in their 60s reported the highest financial losses to fraud, totaling \$1.18 billion in 2024 (AARP, 2025).
Moreover, prior research has shown that online fraud is not only a technological or economic phenomenon but is also shaped by cultural and regional contexts (Parti et al., 2025; Houtti et al., 2024). For instance, a comparative study across 12 countries reveals variations in scam exposure, victimization, types, vectors, and reporting (Houtti et al., 2024). In China, scholars have noted that scammers often exploit local cultural narratives, such as deference to authority and collective values (Li and Ai, 2025), and the online fraud environment differs from Western contexts (Lee, 2021). China is one of the world’s most populous countries and is experiencing rapid population aging (Lobanov-Rostovsky et al., 2023), yet online fraud targeting older adults remains understudied. We therefore situate our study in the Chinese context.
Despite numerous studies in HCI highlighting the unique vulnerabilities of older adults to online fraud and calling for more tailored interventions (Deng et al., 2025; Dai and McGrenere, 2025; Oliveira et al., 2017; Lin et al., 2019), empirical fraud intervention for this population remains scarce. Existing approaches to anti-fraud education for older adults largely rely on traditional one-way formats such as training sessions or broadcast media (Button et al., 2024), which provide limited interactivity and practical exercises. Approaches that are more engaging, such as board games (Chung and Yeung, 2023; Camilleri et al., 2025), have primarily focused on phishing or general cybersecurity rather than online fraud.
@@
To investigate our research questions, we designed three simulation roles, including Experiencer, where older adults engaged in online fraud conversations with an LLM playing the scammer, Helper, where they persuaded an LLM-portrayed victim not to fall into fraud, and Observer, where they watched LLM-generated chat records based on real-world fraud cues. We conducted a between-subjects study with 144 Chinese older adults. To evaluate the intervention, we assessed online fraud awareness, user experience, qualitative feedback, and a follow-up study for retention. Our findings show that role-based simulation was more effective than the control condition in improving fraud cue identification. Both Experiencers and Helpers outperformed the Control group, and Helpers further outperformed Observers. However, these group differences were not observed in the follow-up evaluation. Qualitative feedback informed our reimagination of fraud education design for older adults. Our work makes the following contributions to the human-computer interaction (HCI) and security and privacy (S&P) literature.
@@ 2.1. Online Fraud and Older Adults
Among digital threats, online fraud is one of the most pressing challenges for older adults. The FBI’s Internet Crime Complaint Center documented a 14% increase in elder fraud cases, and tech support scams were the most widely reported kind of elder fraud in 2023 (FBI, 2024). Age-related changes make older adults particularly susceptible to online fraud. Prior research has identified various risk factors, such as cognitive decline (Judges et al., 2017), an overly trusting nature (Shao et al., 2019), and a lack of knowledge about fraud prevention (James et al., 2014). For example, several studies indicated older adults show higher susceptibility to phishing and are more likely to click compared to younger groups (Greitzer et al., 2021; Li et al., 2020; Lin et al., 2019)
@@ 2.3. Fraud Cues
Real-world fraud in China is rarely a single-shot decision problem; it typically unfolds through chat-based interactions in which scammers strategically exploit social interactions on popular messaging platforms (Deng et al., 2025). For instance, scammers often add victims on WeChat (a popular social app in China) and gradually build rapport through repeated promises and performative care (Deng et al., 2025). In such contexts, simply asking older adults to make a broad, binary judgment about whether a case is fraudulent fails to capture the skills they actually rely on when navigating real scam interactions. In contrast, training older adults to identify specific tactics used by scammers could more directly reflect the underlying reasoning processes and offer greater ecological validity and real-world applicability. Building on these insights, in our training materials we define fraud cues as heuristics that users can draw upon to determine whether an interaction might be a scam, and these cues generally map onto deceptive tactics commonly employed by scammers. In this sense, the cues in our study are more similar to the language and content cues in the phishing literature (Molinaro and Bolton, 2018) and focus less on technical details or visual presentation features.
@@ 2.4. Role-based Simulation Learning
Role-based simulation has been widely adopted across diverse educational domains. In medical and nursing training, role-players are extensively used to help students practice clinical and conversational skills in preparation for real patient encounters (Janssens et al., 2023; Wang and Ji, 2021; O’Regan et al., 2016). In computer science, Asanka et al. propose a platform where students actively participate in a role-based learning exercise designed to simulate real-world big data projects (Asanka and Mahanama, 2024). In business contexts, role-play simulations are employed to prepare leaders for interpersonal conflict, negotiations, performance reviews, and ethical dilemmas (Wikipedia, 2025).
Role-based simulation has also seen widespread application in HCI and S&P areas, such as training about system attacks (Cone et al., 2007) and phishing (Karumbaiah et al., 2016; Lain et al., 2022; Gopavaram et al., 2021; Kumaraguru et al., 2008; Yeoh et al., 2022; Chen et al., 2024). For example, CyberCIEGE is a construction and management resource simulation game in which students play the role of decision-makers for enterprises, facing scenarios that require balancing security decisions against organizational needs (Thompson and Irvine, 2014). PhishGuru delivered simulated phishing attacks to users and provided training immediately after participants fell for them (Kumaraguru et al., 2008).
@@ 3.1. Role-playing Learning Modes
Drawn on established learning theories and prior empirical findings, we designed three role-based learning modes that reflect different perspectives on online fraud, shown in Fig. 1.
@@ 3.2. Design Materials
Analysis of our collected cases revealed that although the concrete fraudulent content may differ, for example, false advertising of health products or health courses in the health context, the underlying fraud cues (e.g., exaggerating illness to create anxiety in older adults) being leveraged are largely the same. To extract the online fraud cues, the first author conducted open coding (Corbin and Strauss, 2014) to identify potential fraud cues. Specifically, the author first read through the collected online fraud discourses multiple times and extracted all segments that contained any concrete detail about how the scam was carried out. Based on these extracted segments, the first author developed initial codes that demonstrated the fraud cues. For example, in the segment where “the project manager tells Mr. Qiao that a blockbuster film can earn several times the returns and carries no risk,” we coded this as the fraud cue “promising high returns with no risk.” Through several rounds of team discussions, the co-authors reviewed the extracted segments and codes and worked collaboratively to refine the fraud cues. Finally, we consolidated a set of fraud cues, including 9 cues in the finance context and 7 cues in the health context. [^1] Regarding whether cues within each context are mutually exclusive, the cues of each context in our list do not overlap with one another, but we cannot guarantee mutual exclusivity across all real-world fraud contexts.
We also observed that in a given fraud case, cues rarely appear in isolation. Instead, certain cues frequently co-occur, for example, “promise high returns with low risk”, “claim to be financial experts”, and “pitch senior-living investment schemes”, often appear together in one fraud case. Therefore, to better mirror real-world patterns, our intervention trains older adults using cue sets [^2] A cue from one set may still combine with cues from other sets in real-world cases. Our cue sets only represent combinations commonly observed in our dataset, not the only possible groupings. rather than individual cues. Moreover, our goal is not to have participants learn every cue, but to assess the learning outcomes associated with different perspectives (Experiencer, Helper, Observer). Thus, to avoid overburdening older adults, we only selected 6 cues (i.e., 2 sets) per context. We chose 2 sets in each context so that subsequent evaluation, particularly the fraud cue identification test, would yield more robust results (see Sec. 4.4.1 for details). The selected fraud cues are shown in the Tab. 1, with specific definitions in the Supplementary Material. All these cues were validated by one police officer who is working on the ground for anti-fraud education and three S&P researchers, with more details in Sec. 4.1.
@@
Context Cue Set Cue
@@ B Cue 4: Leverage peer influence
Health A Cue 1: Exaggerating illness to create anxiety
@@
B Cue 4: Framing products as miracle cures with refund guarantees
@@ 3.3.1. Experiencer
Moreover, we provided the set of online fraud cues and their definitions in Tab. 1 to guide the scope of the LLM’s responses. We established strict rules for their use. The model must use, and only use, the predefined online fraud cues and prohibit the introduction of any additional fraud cues to ensure that all cues are systematically covered and remain within our defined boundaries. To maintain the balance and naturalness of these cues, we instructed the model to balance the frequency and depth of each cue, while allowing the order of their appearance to flexibly adapt to the flow of the conversation.
Third, to create a realistic and user-friendly interaction, we also specified the language style. The tone should be natural and realistic, responses were to be concise to avoid distracting older adults, and the total number of conversation turns was controlled to prevent feelings of fatigue or over-engagement. Additionally, we implemented a user-controlled exit option, allowing participants to end the conversation. Finally, we provided several example dialogues for the LLM to reference as guidance. The prompt used in this mode, coupled with few-shot learning (Brown et al., 2020), is provided in the Supplementary Material.
@@ 3.3.2. Helper
GPT-4o served as our backend model to simulate a potential fraud victim. We also first stated our educational goal. Then the LLM was instructed to simulate a potential victim by retelling to the user what the scammer said to it in a turn-taking dialogue, revealing fraud cues for older adults to identify. Moreover, the scammer’s words that the LLM retold were strictly based on the predefined online fraud cues and their definitions in Tab. 1, and the same rules for cue use applied as in the Experiencer mode.
@@ 3.3.3. Observer
In this mode, older adults acted as third-party observers of simulated scam conversations. These conversations were generated using the extracted fraud cues described in Tab. 1. Specifically, we provided the fraud cues as part of the input to GPT-4o and requested it to generate scam dialogues, which were subsequently refined by S&P experts to ensure plausibility. The prompt used for generating these dialogues is provided in the Supplementary Material.
@@ 4. Study Procedure
We conducted a between-subjects experiment to examine the effect of ROLESafe on improving older adults’ online fraud awareness. Fig. 2 shows an overview of the study design.
@@ 4.3. Learning through ROLESafe
    Experiencer group: Participants were told that their task was to chat with another person through a WeChat-style interface. The conversation continued until they no longer wished to proceed. Fig. 3 (a) shows a learning example between an experiencer-group user and the system.
@@
    Helper group: Participants were told that their task was to assist a potential victim of fraud by persuading the individual not to be deceived. They interacted with the victim through the WeChat-style interface, with the conversation lasting as long as they chose. Fig. 3 (b) shows a learning example between a helper-group user and the system.
@@
During the intervention, researchers remained present throughout the session to accompany participants but did not interfere with the tasks. The detailed model performance is shown in Appendix C.
@@ 4.4. Post-Intervention Evaluation
Fraud cue identification test. Fig. 4 shows an example of a fraud cue identification test, including a fraud case and a task. [^5] We did not include non-scam cases because our goal was to assess whether participants could identify targeted fraud cues, rather than to evaluate binary classification performance and adding non-scam cases would shift the construct with limited benefit for the training objective. Participants were presented with an online fraud test case (e.g., Fig. 4 (right)). Then, they needed to complete the cue identification task (e.g., Fig. 4 (left)). The task was “Please select the fraud cues below that you think are present in the above conversation.” Participants selected any number of cues they believed appeared in the case from six possible fraud cues that had been identified for the given context.
@@
Fig. 2 (green parts) illustrates how a participant goes through the learning and evaluation phases. Specifically, after completing the learning phase in one context, participants took a fraud cue identification test corresponding to that context (e.g., finance context). They then proceeded to the second learning phase in the other context and completed another test specific to that context (e.g., health context). Thus, each participant completed two fraud cue identification tests in total. The order of fraud contexts in the learning phases was randomized, and within each context, the tested cue set (A or B) was also randomly assigned. For instance, a participant might work with cue set A in the financial context and cue set B in the health context. This design ensured that participants did not all encounter the same cue set, thereby improving the robustness of the tests.
@@
We adopted a multiple-choice format rather than a per-cue binary response to reduce satisficing behavior observed in the pilot study, where some older participants tended to endorse all cues without careful consideration. The multiple-choice layout encourages comparative thinking and deliberate selection. To evaluate identification accuracy, we used the F1 score, which balances precision and recall, providing a relatively fair measure of overall cue-recognition ability (Powers, 2011). It ranges from 0 to 1, with higher values indicating stronger fraud-cue recognition ability. The final F1 score was calculated as the average of the two tests.
@@
The test cases were generated using the fraud cues collected in Sec. 3.2.3. Specifically, each of the four fraud cue sets from Tab. 1 was provided as part of the input prompts to GPT-4o, which generated four online fraud test cases. Researchers then refined the outputs to ensure plausibility. The prompt used for case generation is provided in the Supplementary Material, and the four complete test cases are provided in the Appendix A.1.

### v2

@@ Abstract.
older adults, online fraud, fraud intervention, anti-fraud education, role-playing, simulation, human-centered security, LLM
journalyear: 2026copyright: ccconference: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems; April 13–17, 2026; Barcelona, Spainbooktitle: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems (CHI ’26), April 13–17, 2026, Barcelona, Spaindoi: 10.1145/3772318.3791003isbn: 979-8-4007-2278-3/2026/04ccs: Security and privacy Human and societal aspects of security and privacyccs: Human-centered computing Human computer interaction (HCI)
@@ 1. Introduction
Online fraud has emerged as one of the most prevalent and costly digital threats worldwide (Button et al., 2014; Cross, 2022). Older adults are disproportionately targeted due to multiple age-related risk factors (Judges et al., 2017; Alves and Wilson, 2008; James et al., 2014). Their vulnerability is further exacerbated by the continuous evolution and increasing sophistication of scam tactics (Deng et al., 2025; Zhai et al., 2025) and AI’s application in enhancing the effectiveness, speed, and credibility of fraudulent schemes (Herrera et al., 2024). Consequently, older adults often express growing apprehension about the complexity of such threats and the difficulty of recognizing fraud in practice (Sheil et al., 2024). Fraud targeting older adults has increasingly become a global threat. In China, survey data from the Chinese Academy of Social Sciences revealed that 17.25% of older users had experienced online fraud in 2022 (Neican, 2022). According to the U.S. Federal Trade Commission, American victims in their 60s reported the highest financial losses to fraud, totaling \$1.18 billion in 2024 (AARP, 2025).
Moreover, online fraud is not only a technological or economic phenomenon but is also shaped by cultural and regional contexts (Parti et al., 2025; Houtti et al., 2024). For instance, a comparative study across 12 countries reveals variations in scam exposure, victimization, types, vectors, and reporting (Houtti et al., 2024). In China, where the online fraud environment differs from Western contexts (Lee, 2021), scholars have noted that scammers often exploit local cultural narratives, such as deference to authority and collective values (Li and Ai, 2025). China is one of the world’s most populous countries and is experiencing rapid population aging (Lobanov-Rostovsky et al., 2023), yet online fraud targeting older adults remains understudied. We therefore situate our study in the Chinese context.
Despite numerous studies in HCI highlighting the unique vulnerabilities of older adults to online fraud and calling for more tailored interventions (Deng et al., 2025; Dai and McGrenere, 2025; Oliveira et al., 2017; Lin et al., 2019), empirical fraud intervention for this population remains scarce. Existing approaches to anti-fraud education for older adults largely rely on traditional one-way formats such as training sessions or broadcast media (Button et al., 2024), which provide limited interactivity and practical exercises. More engaging approaches, such as board games (Chung and Yeung, 2023; Camilleri et al., 2025), have primarily focused on phishing or general cybersecurity rather than online fraud.
@@
To investigate our research questions, we designed three simulation roles, including Experiencer, where older adults engaged in online fraud conversations with an LLM playing the scammer, Helper, where they persuaded an LLM-portrayed victim not to fall into fraud, and Observer, where they watched LLM-generated chat records based on real-world fraud cues. We conducted a between-subjects experiment with 144 Chinese older adults and evaluated our intervention’s effectiveness in raising online fraud awareness and user experience, while gathering participants’ qualitative feedback. Our findings show that role-based simulation was more effective than the control condition in improving fraud cue identification. Both Experiencers and Helpers outperformed the Control group, and Helpers further outperformed Observers. However, these group differences were not observed in a follow-up evaluation. Qualitative feedback informed our reimagination of fraud education design for older adults. Our work makes the following contributions to the human-computer interaction (HCI) and security and privacy (S&P) literature.
@@ 2.1. Online Fraud and Older Adults
Online fraud is one of the most pressing digital threats for older adults. The FBI’s Internet Crime Complaint Center documented a 14% increase in elder fraud cases, and tech support scams were the most widely reported kind of elder fraud in 2023 (FBI, 2024). Age-related changes make older adults particularly susceptible to online fraud. Prior research has identified various risk factors, such as cognitive decline (Judges et al., 2017), an overly trusting nature (Shao et al., 2019), and a lack of knowledge about fraud prevention (James et al., 2014). For example, several studies indicated older adults show higher susceptibility to phishing and are more likely to click compared to younger groups (Greitzer et al., 2021; Li et al., 2020; Lin et al., 2019).
@@ 2.3. Fraud Cues
Real-world fraud in China is rarely a single-shot decision problem; it typically unfolds through chat-based interactions in which scammers strategically exploit social interactions on popular messaging platforms (Deng et al., 2025). For instance, scammers often add victims on WeChat (a popular social app in China) and gradually build rapport through repeated promises and performative care (Deng et al., 2025). In such contexts, simply asking older adults to make a broad, binary judgment about whether a case is fraudulent fails to capture the skills they actually rely on when navigating real scam interactions. In contrast, training older adults to identify specific tactics used by scammers could more directly reflect the underlying reasoning processes and offer greater ecological validity and real-world applicability. Building on these insights, in our training materials, we define fraud cues as heuristics that users can draw upon to determine whether an interaction might be a scam, and these cues generally map onto deceptive tactics commonly employed by scammers. In this sense, the cues in our study are more similar to the language and content cues in the phishing literature (Molinaro and Bolton, 2018) and focus less on technical details or visual presentation features.
@@ 2.4. Role-based Simulation Learning
Role-based simulation has been widely adopted across diverse educational domains. In medical and nursing training, role-players are extensively used to help students practice clinical and conversational skills before encountering real-world patients (Janssens et al., 2023; Wang and Ji, 2021; O’Regan et al., 2016). In computer science, Asanka et al. propose a platform where students actively participate in a role-based learning exercise designed to simulate real-world big data projects (Asanka and Mahanama, 2024). In business contexts, role-play simulations are employed to prepare leaders for interpersonal conflict, negotiations, and performance reviews (Wikipedia, 2025).
Role-based simulation has also seen widespread application in HCI and S&P areas, such as training about system attacks (Cone et al., 2007) and phishing (Karumbaiah et al., 2016; Lain et al., 2022; Gopavaram et al., 2021; Kumaraguru et al., 2008; Yeoh et al., 2022; Chen et al., 2024). For example, CyberCIEGE is a construction and management resource simulation game in which students act as decision-makers for enterprises, facing scenarios that require balancing security decisions against organizational needs (Thompson and Irvine, 2014). PhishGuru delivered simulated phishing attacks and provided training immediately after participants fell for them (Kumaraguru et al., 2008).
@@ 3.1. Role-playing Learning Modes
Drawn on established learning theories and prior empirical findings, we designed three role-based learning modes that reflect different perspectives on online fraud, shown in Figure 1.
@@ 3.2. Design Materials
Our analysis of collected cases revealed that although the concrete fraudulent content may differ—for example, health-related fraud could have false advertising of health products or courses—the underlying fraud cues leveraged, such as exaggerating illness to create anxiety in older adults, are largely the same. To extract the online fraud cues, the first author conducted open coding (Corbin and Strauss, 2014) to identify potential fraud cues. Specifically, the author first read through the collected online fraud discourses multiple times and extracted all segments that contained any concrete detail about how the scam was carried out. Based on these extracted segments, the first author developed initial codes that demonstrated the fraud cues. For example, in the segment where “the project manager tells Mr. Qiao that a blockbuster film can earn several times the returns and carries no risk,” we coded this as the fraud cue “promising high returns with no risk.” Through several rounds of team discussions, the co-authors reviewed the extracted segments and codes and worked collaboratively to refine the fraud cues. Finally, we consolidated a set of fraud cues, including nine cues in the finance context and seven cues in the health context. [^1] Regarding whether cues within each context are mutually exclusive, the cues of each context in our list do not overlap with one another, but we cannot guarantee mutual exclusivity across all real-world fraud contexts.
We also observed that in a given fraud case, cues rarely appear in isolation. Instead, certain cues frequently co-occur, for example, “promise high returns with low risk,” “claim to be financial experts,” and “pitch senior-living investment schemes” often appear together in one fraud case. Therefore, to better mirror real-world patterns, our intervention trains older adults using cue sets [^2] A cue from one set may still combine with cues from other sets in real-world cases. Our cue sets only represent combinations commonly observed in our dataset, not the only possible groupings. rather than individual cues. Moreover, our goal is not to have participants learn every cue, but to assess the learning outcomes associated with different perspectives (Experiencer, Helper, Observer). Thus, to avoid overburdening older adults, we only selected six cues (i.e., two sets) per context. We chose two sets in each context so that subsequent evaluation, particularly the fraud cue identification test, would yield more robust results (see Sec. 4.4.1 for details). The selected fraud cues are shown in Table 1, with full fraud cues and their specific definitions in the Supplementary Material. All these cues were validated by one police officer who is working on the ground for anti-fraud education and three S&P researchers, with more details in Sec. 4.1.
@@
Table 1: The online fraud cues within finance and health contexts.
Context Set Cue
@@ B Cue 4: Leverage peer influence
Health A Cue 1: Exaggerate illness to create anxiety
@@
B Cue 4: Frame products as miracle cures with refund guarantees
@@ 3.3. Learning Mode Implementation
We chose LLM over other methods (e.g., Wizard-of-Oz) to implement these learning modes for both practical and conceptual reasons. Our long-term goal is to develop scalable tools for online fraud education, and using an LLM allows us to test an intervention format closer to real deployment. The study also involved multi-turn, role-based interactions across several conditions, making a human-wizard approach labor-intensive and prone to inconsistency and experimenter bias. This study serves as an initial demonstration of the feasibility of an LLM-based intervention, with future work needed to further refine dialogue design.
@@ 3.3.1. Experiencer
Moreover, we provided the set of online fraud cues and their definitions in Table 1 to guide the scope of the LLM’s responses. We established strict rules in this implementation. The model must use, and only use, the predefined online fraud cues and prohibit the introduction of any additional fraud cues to ensure that all cues are systematically covered and remain within our defined boundaries. To maintain the balance and naturalness of these cues, we instructed the model to balance the frequency and depth of each cue, while allowing the order of their appearance to flexibly adapt to the flow of the conversation.
Third, to create a realistic and user-friendly interaction, we specified the language style. The tone should be natural and realistic, responses were to be concise to avoid distracting older adults, and the total number of conversation turns was controlled to prevent feelings of fatigue or over-engagement. Additionally, we implemented a user-controlled exit option, allowing participants to end the conversation. Finally, we provided several example dialogues for the LLM to reference as guidance. The prompt used in this mode, coupled with few-shot learning (Brown et al., 2020), is provided in the Supplementary Material.
@@ 3.3.2. Helper
GPT-4o served as our backend model to simulate a potential fraud victim. We also stated our educational goal. Then the LLM was instructed to simulate a potential victim by retelling to the user what the scammer said to it in a turn-taking dialogue, revealing fraud cues for older adults to identify. Moreover, the scammer’s words that the LLM retold were strictly based on the predefined online fraud cues and their definitions in Table 1, and the same rules for cue use applied as in the Experiencer mode.
@@ 3.3.3. Observer
In this mode, older adults acted as third-party observers of simulated scam conversations. These conversations were generated using the extracted fraud cues described in Table 1. Specifically, we provided the fraud cues as part of the input to GPT-4o and requested it to generate scam dialogues, which were subsequently refined by S&P experts to ensure plausibility. The prompt used for generating these dialogues is provided in the Supplementary Material.
Figure 2: Study overview. For the study design, participants were recruited and assigned to one of four groups: control, experiencer, helper, or observer. Each group engaged in different learning activities, such as reading materials, experiencing or observing scams, or helping a victim. Post-intervention evaluation included fraud awareness assessment, user experience evaluation, and feedback interviews, followed by a one-month fraud awareness follow-up assessment. For the study procedure, each participant was first assigned to a group. They then learned the fraud cues in one context (e.g., finance A) and completed a fraud-cue identification test in the same context. Next, they learned the fraud cues in the other context (e.g., health B) and completed the corresponding identification test, followed by a self-efficacy scale, user-experience evaluation, and feedback interview. About one month later, the participant completed fraud-cue identification tests again in both contexts (e.g., finance A and health B) and repeated the self-efficacy scale.
@@ 4. Study Procedure
We conducted a between-subjects experiment to examine the effect of ROLESafe on improving older adults’ online fraud awareness. Figure 2 shows an overview of the study design.
@@ 4.3. Learning through ROLESafe
Figure 3: Examples of how users interacted with the LLM during the learning phase in the finance-B context. In the experiencer condition (left), the model plays the role of a scammer attempting to persuade the user, revealing fraud cues such as leveraging peer influence, invoking fabricated government-backed investment schemes, and promising shareholder dividends. In contrast, in the helper condition (right), the model adopts a supportive stance, guiding the user to question suspicious claims and recognize common fraud tactics.
@@
    Experiencer group: Participants were told that their task was to chat with another person through a WeChat-style interface. The conversation continued until they no longer wished to proceed. Figure 3a shows a learning example between an experiencer-group user and the system.
@@
    Helper group: Participants were told that their task was to assist a potential victim of fraud by persuading the individual not to be deceived. They interacted with the victim through the WeChat-style interface, with the conversation lasting as long as they chose. Figure 3b shows a learning example between a helper-group user and the system.
@@
During the intervention, researchers remained present throughout the session to accompany participants but did not interfere with the tasks. The detailed model performance is shown in Sec. 6.4.
@@ 4.4. Post-Intervention Evaluation
Figure 4: Fraud-cue identification task used in the finance context. Participants were shown a conversational fraud case (right) and asked to select which cues were present from a predefined list (left). The highlighted segments in the conversation map onto specific fraud cues, such as leveraging peer influence, invoking fabricated government-backed projects, and promising shareholder dividends, demonstrating how these cues were embedded in the scenario and surfaced for recognition during the test.
@@
Fraud cue identification test. Figure 4 shows an example of a fraud cue identification test, including a fraud case and a task. [^5] We did not include non-scam cases because our goal was to assess whether participants could identify targeted fraud cues, rather than to evaluate binary classification performance and adding non-scam cases would shift the construct with limited benefit for the training objective. Participants were presented with an online fraud test case before completing the cue identification task. The task was “Please select the fraud cues below that you think are present in the above conversation.” Participants selected any number of cues they believed appeared in the case from six possible fraud cues that had been identified for the given context.
@@
Figure 2 (green parts) illustrates how a participant goes through the learning and evaluation phases. Specifically, after completing the learning phase in one context, participants took a fraud cue identification test corresponding to that context (e.g., finance context). They then proceeded to the second learning phase in the other context and completed another test specific to that context (e.g., health context). Thus, each participant completed two fraud cue identification tests in total. The order of fraud contexts in the learning phases was randomized, and within each context, the tested cue set (A or B) was also randomly assigned. For instance, a participant might work with cue set A in the financial context and cue set B in the health context. This design ensured that participants did not all encounter the same cue set, thereby improving the robustness of the tests.
@@
We adopted a multiple-choice format rather than a per-cue binary response to reduce satisficing behavior observed in the pilot study, where some older participants tended to endorse all cues without careful consideration. The multiple-choice layout encourages comparative thinking and deliberate selection. To evaluate identification accuracy, we used the F1 score, which balances precision and recall, providing a relatively fair measure of overall cue-recognition ability (Powers, 2011). We computed the F1 score by taking twice the number of options that were selected in both the predicted and ground-truth answers, and dividing it by the total number of options selected across the predicted and the ground-truth answers. It ranges from 0 to 1, with higher values indicating stronger fraud-cue recognition ability. The final F1 score was calculated as the average of the two tests.
@@
The test cases were generated using the fraud cues collected in Sec. 3.2.3. Specifically, each of the four fraud cue sets from Table 1 was provided as part of the input prompts to GPT-4o, which generated four online fraud test cases. Researchers then refined the outputs to ensure plausibility. The prompt used for case generation is provided in the Supplementary Material, and the four complete test cases are provided in the Appendix A.1.